OpenStudy (anonymous):

Classify the solution as acidic, basic, neutral [OH^-]=9.8 X 10^-11M

OpenStudy (anonymous):

\[[\text{H}^{1+}] = \frac{k_w}{[\text{OH}^{1-}]}, \qquad k_w = 1.0 \times 10^{-14} \quad \text{at SATP}\] Use this to determine the concentration of the hydrogen (or hydronium) ion. The solution will be acidic if the concentration of the hydrogen ion exceeds \[1 \times 10^{-7}~\text{M}\]